Consider the following problem involving three bonds that pay coupons annually. All three bonds have face values of $1,000. Bond A pays annual coupons of $100, and is worth $1,075. Bond B pays annual coupons of $50, and is worth $985. The term structure of interest rates is not flat.
2. What is the value of Bond C, if it pays annual coupons of $150?
Answer #1

From Bond A:
PV of 100+PV of 1000=1075
=>PV of 1000=1075-PV of 100

From Bond B:
PV of 50+PV of 1000=985
PV of 50=985-PV of 1000
PV of 50=985-(1075-PV of 100)
=>PV of 50-PV of 100=-90
=>PV of 100-PV of 50=90

PV of 100=2*PV of 50
So, PV of 50=90
PV of 100=2*PV of 50=2*90=180

PV of 1000=1075-180=895

Bond C Price=PV of 150+PV of 1000=3*PV of 50+PV of 1000=3*90+895=$1,165.00